# Planet formation theory: an overview

### v1

@@ 3 Aerodynamic and collisional evolution of solids
Dust inherited by the disk is assumed to follow the size distribution, $n(a)\propto a^{-3.5}$, that is inferred from extinction measurements in the interstellar medium (Mathis et al. 1977). Growth up to the scale of small macroscopic solids (“pebbles”, which might be anywhere between mm and meters in physical size) is controlled by aerodynamic and collisional processes. The physics of these processes is quite subtle, but most of the important results follow from two considerations:
@@
    Solid particles are coupled to the gas via aerodynamic drag forces, but imperfectly so. Deviations from perfect coupling have multiple consequences, including vertical settling, radial drift, particle concentration in vortices, and small-scale clustering and collisions. Back-reaction of the aerodynamic forces on the gas, when the dust to gas density ratio is non-negligible, leads to the streaming instability. These effects depend not on the physical particle size, but on a dimensionless measure of “size” for aerodynamic purposes called the Stokes number.
@@
Figure 2: Aerodynamic coupling between solid particles and disk gas controls the early phases of planet formation on multiple scales. On the smallest scale, it leads to clustering of solids (simulation from Chan et al. 2024) and collisions between individual particles (Ormel and Cuzzi 2007). On intermediate scales, it leads to the streaming instability (Youdin and Goodman 2005) and planetesimal formation (simulation from Simon et al. 2016). The Kuiper Belt contact binary Arrokoth (bottom right panel) is consistent with this formation channel. On large, observable scales, aerodynamics leads to vertical settling, radial drift, and particle concentration at pressure maxima.
@@
Small dust grains in gaseous disks experience low velocity collisions due to Brownian motion (Dullemond and Dominik 2005), that lead to sticking and the formation of porous aggregates. Larger particles collide at higher velocities, due to coupling to small-scale turbulence or size-dependent drift, leading to compactification and, eventually, bouncing or fragmentation. A microscopic theory (“JKRS theory”, applied to the astrophysical context by Dominik and Tielens 1997) relates the collision outcome to the surface energies of the materials involved, and can be used to model the growth of aggregates via molecular dynamics techniques (e.g. Wada et al. 2009). The fiducial results for grain growth, however, come from laboratory experiments, usually performed in microgravity conditions (Blum and Wurm 2008). For silicate particles, considered to be representative of the inner disk interior to the snowline, net growth occurs for collisions with relative velocity $\Delta v<1\ {\rm ms}^{-1}$ (Blum and Wurm 2008). For water ice, collision experiments by Gundlach and Blum (2015) at $T\approx 100-200\ {\rm K}$ found growth for $\Delta v<10\ {\rm ms}^{-1}$. Measurements of water ice surface energies, however, are similar to values measured for silicates at temperatures $T<175\ {\rm K}$ (Musiolik and Wurm 2019). In JKRS theory this would imply that both silicate and ice aggregates cease to stick above similar (low) threshold collision speeds, except near the snowline where water ice unambiguously shows a higher threshold. There is evidently some uncertainty in this key quantity, and it is common in models to consider the sticking and fragmentation thresholds for ice as parameters and explore the effect of different values in the 1-10 ms-1 range.
@@
Translating the experimental results for the sticking and fragmentation thresholds into a predicted maximum particle size $a_{\rm max}$ requires knowledge of what sets $\Delta v$ as a function of size $a$. In the most common situation, collisions at $a\sim a_{\rm max}$ are caused by aerodynamic coupling to small-scale turbulence (this process also leads to small-scale clustering of solids; Cuzzi et al. 2001). For particles of size $a_{1}$ and $a_{2}$, coupled aerodynamically to isotropic Kolmogorov turbulence, Ormel and Cuzzi (2007) provide the standard (in the astrophysical literature) calculation that gives $\Delta v(a_{1},a_{2})$. Adopting an $\alpha$ model for the turbulence, the peak collision speed is found to be,
@@
This allows observations of the thickness of dust disks to be used to constrain disk turbulence. The assumption that $D=\nu$ is quite crude, and there can be substantial deviations from this simple formula in disks where the turbulence is driven by the Vertical Shear Instability (Stoll and Kley 2016; Pfeil et al. 2024; Fukuhara and Okuzumi 2024) or by self-gravity (Baehr and Zhu 2021).
@@
Even though $\eta\sim(h/r)^{2}$ is small, typically only $\sim{\rm few}\times 10^{-3}$, the rate of the resulting drift is highly consequential. At 1 au, the time scale for drift of particles with $\tau_{\rm S}=1$ would be as short as of the order of a thousand years. Particles with this Stokes number would be of the order of meters in size, leading to the moniker “meter-sized barrier” for this aerodynamic constraint on planetesimal formation. (Material physics limits on sticking already make the growth of meter-sized boulders improbable, but the meter-sized barrier provides an additional, independent, argument against planetesimal formation by pairwise coagulation.) More importantly in a modern context, particles with physical sizes of mm-cm and Stokes numbers $\tau_{\rm S}\sim 10^{-2}$, that can readily form given our best understanding of coagulation physics, are predicted to suffer substantial drift even in the outer disk on scales of 10-100 au.
@@
Turbulent diffusion in the radial direction can be important both for the evolution of trace gas species, and for the evolution of dust populations that are simultaneously subject to radial drift (Morfill and Voelk 1984; Clarke and Pringle 1988). Tominaga et al. (2019) derive formulae for dust diffusion in disks that conserve angular momentum and which can be used to model dust concentration processes and possible instabilities. An important consequence of diffusion is the qualitative change it makes to how dust evolves in the vicinity of gas pressure maxima. Ignoring diffusion, an axisymmetric pressure maximum – whether it forms spontaneously or at a planetary gas edge – is efficient at trapping aerodynamically drifting solids, filtering all but the smallest particles that are swept inward by the mean flow (Rice et al. 2006). With diffusion, it is always possible to find a steady state solution in which the turbulent flux overcomes the aerodynamic trapping effect, such that pressure maxima temporarily detain but do not permanently imprison solids (Ward 2009; Zhu et al. 2012b).
@@ 4 Planetesimal formation
If $\rho_{\rm dust}>\rho_{\rm Roche}$, the solid material can collapse gravitationally to form planetesimals. Although there are circumstances where this process may operate, it is no longer considered to be a generally viable planetesimal formation mechanism. Typically, vertical shear will generate enough turbulence to halt settling before the critical density is reached, even in an intrinsically non-turbulent disk (Cuzzi et al. 1993). The other obvious mechanism for planetesimal formation is a continuation of pairwise sticking collisions from pebble to planetesimal scales (Weidenschilling and Cuzzi 1993). Again, while this model cannot be entirely ruled out, it is largely unviable given current understanding of fragmentation and radial drift processes.
The current working hypothesis for planetesimal formation is that the streaming instability (Youdin and Goodman 2005) generates particle clumps that are dense enough to collapse gravitationally to planetesimals. The streaming instability is a local, linear instability of the Nakagawa et al. (1986) solution for equilibrium radial drift. It requires consideration of the momentum feedback particles exert on the gas, and a non-zero drift velocity (hence, it cannot formally be operative at the peak of a pressure maximum). Youdin and Goodman (2005) show that the streaming instability has non-zero (though sometimes small) growth rates for almost all combinations of Stokes number and dust to gas ratio, and it should therefore be generically present in protoplanetary disks. Jacquet et al. (2011) and Magnan et al. (2024) discuss the rather subtle physical mechanism of the streaming instability, which is the most important (for planet formation) of a broader class of resonant drag instabilities (Squire and Hopkins 2018). It should be noted that the Nakagawa et al. (1986) equilibrium studied by Youdin and Goodman (2005) differs in multiple ways from both real disks and from the setups for most streaming instability simulations. The Nakagawa et al. (1986) model is vertically unstratified (vertical gravity is generally not negligible), does not include turbulent diffusion (which is present if there is any instrinsic disk turbulence and which is needed to establish a stratified particle layer equilibrium), and is based on a fluid description of the particles (this is only a good approximation for $\tau_{\rm S}\ll 1$, and is not used in particle-based simulations). Subsequent linear calculations have included the effects of a particle size distribution (Krapp et al. 2019; Paardekooper et al. 2020; Zhu and Yang 2021) and turbulent diffusion (Chen and Lin 2020; Umurhan et al. 2020; Gerbig et al. 2024), which act to reduce the growth rates and restrict the range of parameters over which linear instability is present.
@@
A second important simulation result is the predicted mass function of planetesimals formed from the streaming instability. A large number of studies agree that streaming-initiated collapse leads to a top-heavy initial planetesimal mass function, in which most of the total mass is in the largest objects (Johansen et al. 2007; Simon et al. 2016; Schäfer et al. 2017; Li et al. 2019; Schäfer et al. 2024). The absolute size of the largest planetesimals is of course a function of the disk conditions, but for reasonable parameters it is often hundreds of kilometers. This first principles prediction is large compared to what was assumed in most formation-agnostic planetary growth models in the past, and has substantial downstream implications both for the Solar System (“asteroids were born big”; Morbidelli et al. 2009) and for planet formation generally. In particular, 100+ km planetesimals experience much less aerodynamic drag than 100 m-1 km planetesimals, and when their orbits are excited by gravitational perturbations the population will saturate at a higher velocity dispersion. This reduces the rate at which protoplanets grow by planetesimal accretion, so that, counter-intuituvely, being able to rapidly make large primordial planetesimals does not necessarily imply faster overall planet formation time scales.
There is no analytic expectation for the form of the initial planetesimal mass function, but several simple functions (an exponentially tapered power-law, a broken power-law, etc) provide fits to numerical results that can be compared to Solar System populations of small bodies. The Kuiper Belt is widely considered to be the most primordial surviving population and offers the greatest potential for a test, though even there careful consideration has to be given to the extent of collisional evolution (Morbidelli et al. 2021; Benavidez et al. 2022). Other possible tests of the predictions of planetesimal formation models include the inclination distribution of Kuiper Belt binaries (Nesvorný et al. 2019), and the detailed morphology of the single small Kuiper Belt Object imaged by the New Horizons mission (Arrokoth; McKinnon et al. 2020). The results of these tests are consistent with streaming instability simulations.
@@ 5 Formation of rocky planets, cores, and giants
    Planetesimal accretion. The plantesimals physically collide, due to gravitational interactions, at low enough velocity to grow. The rate of collisions is determined, through the physics of gravitational focusing, by the velocity dispersion of the planetesimals. To a first approximation, one imagines the protoplanet growing from a local planetesimal reservoir that does not drift radially.
@@
    Pebble accretion. While the gas disk is still present, a fraction of appropriately aerodynamically coupled pebbles (typically, $10^{-3}\leq\tau_{\rm S}\leq 1$) that enter the protoplanet’s Hill sphere are gravitationally captured and accreted via a settling flow. The protoplanet accretes a fraction of the solids that drift past its orbit, with an efficiency that depends on its mass, on the Stokes number, and on the disk properties.
@@
Applying the condition that the settling time scale $t_{\rm settle}=b/v_{\rm settle}$ equals the encounter time scale $t_{\rm enc}=b/\delta v$ we find the relation that $\delta v=GMt_{\rm stop}/b^{2}$, and the pebble accretion rate is,
@@
Pebble accretion has several distinctive differences from planetesimal accretion. There is no regime of pebble accretion that exhibits runaway growth. It is also fairly strictly bounded in planet mass. Pebble accretion is inefficient below a pebble initiation mass (Visser and Ormel 2016), and for typical parameters would only be able to directly grow planetesimals if their initial sizes are rather large (at least several hundred km). Pebble accretion is also likely to be inefficient above a pebble isolation mass (Lambrechts et al. 2014; Bitsch et al. 2018), typically of the order of $50M_{\oplus}$. Above this mass, a growing planet starts to open a gap in the disk, forming a local pressure maximum that can halt the radial drift of solids.
@@
Planetesimal accretion is severely limited at large orbital radii by the tendency of more massive bodies to scatter, rather than accrete, planetesimals (this is the origin of the long-known difficulty in forming Uranus and Neptune in situ from planetesimals; Levison and Stewart 2001). Pebble accretion, conversely, tends to work better at moderately large orbital radii – solids at the drift or fragmentation limit have more favorable aerodynamic properties, and there is no reduction in accretion efficiency as the ratio of Hill sphere size to planetary radius increases. Outside perhaps 5-10 au, pebble accretion is presumptively the dominant mechanism for planetary growth (e.g. Drazkowska et al. 2023). More general statements as to the relative role of planetesimals and pebbles in planet formation are, however, hazardous. Planetesimal and pebble accretion are distinct processes that cannot be easily compared, and while pebble accretion rates can be very high this is, in part, a consequence of the rapid radial drift of solids which cannot be sustained indefinitely. Which process dominates in the formation of observed planetary systems can only be considered in the context of a global theoretical model for planet formation (e.g. Bitsch et al. 2015), or by appealing to observational cosmochemical constraints that are only available for the Solar System. Johansen et al. (2021) argue for a large pebble contribution in terrestrial planet formation, but the division between the two processes in a hybrid Solar System model remains the subject of vigorous debate.
@@
Accretion of the primary atmospheres or envelopes of massive planets occurs (in the consensus view) via the process of core accretion (Mizuno 1980), an umbrella term that encompasses the classical planetesimal-only model for Jupiter and Saturn’s formation (Pollack et al. 1996), along with more recent models that can accommodate the existence of the dilute core inferred for Jupiter from JUNO data (Wahl et al. 2017; Helled et al. 2022), or a primary role for pebble accretion (Lambrechts et al. 2014). The key insight of core accretion is the predicted existence of a sharp increase in the gas accretion rate when the envelope mass $M_{\rm env}$ exceeds a critical mass, that is approximately set at $M_{\rm env}\sim M_{\rm core}$ by the onset of self-gravity in the envelope (Rafikov 2006). Below the critical mass, low mass gaseous protoplanets grow relatively slowly, at a rate that is determined by the growth of the core through pebble and planetesimal accretion, and by the time scale on which the envelope can lose thermal energy and contract (Ikoma et al. 2000). If the gas disk dissipates during this initial phase of core accretion, the expected outcome is a planet that may resemble the Solar System’s ice giants, or extrasolar sub-Neptunes. If, instead, the critical mass can be exceeded during the disk lifetime, the planet enters a phase of accelerated gas accretion (this is known as “runaway growth”, although it is still a quasi-hydrostatic evolution) and the outcome is a gas giant with $M_{\rm env}\gg M_{\rm core}$. The minimum core mass that can lead to gas giant formation is primarily a function of the opacity in the envelope, and is roughly $5-10\ M_{\oplus}$ (Piso et al. 2015) for grain size distributions that arise in coagulation-fragmentation equilibrium calculations. Smaller critical core values are possible if envelopes are effectively grain-free (Hori and Ikoma 2010). In the other direction, substantially more massive cores can avoid entering runaway given sufficient heating from planetesimal accretion.
@@
Following runaway, growth of giant planets to Jovian or super-Jovian masses is ultimately limited by the ability of the planet to accrete from across the gap that opens within the protoplanetary disk. Circumplanetary disks, fed by out-of-plane flows from the protoplanetary disk (Tanigawa et al. 2012), may form during this final phase of accretion. Circumplanetary disk formation requires not only a sufficiently large planet mass relative to the thermal mass (roughly, a mass such that $r_{\rm Hill}>h$; e.g. Sagynbayeva et al. 2024), but also favorable thermodynamic conditions that allow fast enough cooling on sub-Hill scales (Ayliffe and Bate 2009; Szulágyi et al. 2016; Fung et al. 2019). Multi-fluid radiation hydrodynamic simulations suggest that cooling on a time scale that is an order of magnitude shorter than the orbital time scale is necessary for gas to form a near-Keplerian disk on scales of $r_{\rm Hill}/3$ (Krapp et al. 2024). Observationally, the detection and characterization of circumplanetary disks is in its infancy. PDS 70c, a young massive exoplanet in a system whose protoplanetary disk has a transition disk morphology, shows sub-mm emission at the location of the planet that is consistent with circumplanetary disk expectations (Isella et al. 2019; Benisty et al. 2021). As illustrated in Figure 4, continuum emission is just one of many potential tracers of circumplanetary disks, which may also be observable in molecular (kinematic) or atomic line emission, or even through the production of jets (Gressel et al. 2013; Zhu 2015).
The consensus view is that the Solar System’s planets, and at least the vast majority of extrasolar planets, formed via core accretion (which, to be clear, is a broad framework for planet formation rather than a precisely specified model). It is possible that some giant planets, especially those approaching the $13\ M_{J}$ threshold that conventionally demarcates planets from brown dwarfs, instead formed via disk instability (Boss 1997). The formation of protostellar disks may well lead to initially self-gravitating disks (Lin and Pringle 1990; Xu and Kunz 2021), which have been observed to fragment into multiple stellar systems (Tobin et al. 2016). The physics of self-gravitating disks is reviewed by Kratter and Lodato (2016). Such disks fragment if the cooling time scale is sufficiently short (of the order of the orbital time scale; Gammie 2001; Rice et al. 2005), conditions that occur only at large ($\sim 50$ au) radii (Clarke 2009; Rafikov 2009). What are currently the most realistic radiation hydrodynamics simulations, by Xu et al. (2024), find that the distribution of initial fragment masses in this regime peaks at about $20\ M_{J}$. Further accretion in what remains a gas-rich environment post-fragmentation would likely increase fragment masses further (Kratter et al. 2010b). (This fate could only be averted if fragments migrate inward rapidly and are partially tidally stripped; Nayakshin 2010; Baruteau et al. 2011; Zhu et al. 2012a). Protostellar disks can also be driven to fragment by sufficiently rapid infall (Kratter et al. 2010a), but this also leads to characteristic masses that are typically in the brown dwarf or low-mass star regime.

### v2

@@ 3 Aerodynamic and collisional evolution of solids
Dust inherited by the disk is expected to follow the size distribution, $n(a)\propto a^{-3.5}$, that is inferred from extinction measurements in the interstellar medium (Mathis et al. 1977). Growth up to the scale of small macroscopic solids (“pebbles”, which might be anywhere between mm and meters in physical size) is controlled by aerodynamic and collisional processes. The physics of these processes, reviewed by Birnstiel (2024), is quite subtle, but most of the important results follow from two considerations:
@@
    Solid particles are coupled to the gas via aerodynamic drag forces, but imperfectly so. Deviations from perfect coupling have multiple consequences, including vertical settling, radial drift, particle concentration in vortices, and small-scale clustering and collisions. Back-reaction of the aerodynamic forces on the gas, when the dust to gas density ratio is non-negligible, leads to additional effects, including most notably the streaming instability. These effects depend not on the physical particle size, but on a dimensionless measure of “size” for aerodynamic purposes called the Stokes number (equation 9).
@@
Figure 2: Aerodynamic coupling between solid particles and disk gas controls the early phases of planet formation on multiple scales. On the smallest scale, it leads to clustering of solids (simulation from Chan et al. 2024). The characteristic clustering scale is the Kolmogorov scale defined for isotropic fluid turbulence – of the order of km at 5 au (Pan et al. 2011). It also results in collisions between individual particles (Ormel and Cuzzi 2007). On intermediate scales, it leads to the streaming instability (Youdin and Goodman 2005) and planetesimal formation (simulation in a $(0.2h)^{3}$ domain from Simon et al. 2016). The Kuiper Belt contact binary Arrokoth (bottom right panel) is consistent with this formation channel. On large, observable scales, aerodynamics leads to vertical settling, radial drift, and particle concentration at pressure maxima.
@@
Small dust grains in gaseous disks experience low velocity collisions due to Brownian motion (Dullemond and Dominik 2005), that lead to sticking and the formation of porous aggregates. Larger particles collide at higher velocities, due to coupling to small-scale turbulence or size-dependent drift, leading to compactification and, eventually, bouncing or fragmentation. A microscopic theory (“JKRS theory”, applied to the astrophysical context by Dominik and Tielens 1997) relates the collision outcome to the surface energies of the materials involved, and this theory can be used to model the growth of aggregates via molecular dynamics techniques (e.g. Wada et al. 2009). The fiducial results for grain growth, however, come from laboratory experiments, usually performed in microgravity conditions (Blum and Wurm 2008). For silicate particles, considered to be representative of the inner disk interior to the snowline, net growth occurs for collisions with relative velocity $\Delta v<1\ {\rm ms}^{-1}$ (Blum and Wurm 2008). For water ice, collision experiments by Gundlach and Blum (2015) at $T\approx 100-200\ {\rm K}$ found growth for $\Delta v<10\ {\rm ms}^{-1}$. Measurements of water ice surface energies, however, are similar to values measured for silicates at temperatures $T<175\ {\rm K}$ (Musiolik and Wurm 2019). In JKRS theory this would imply that both silicate and ice aggregates cease to stick above similar (low) threshold collision speeds, except near the snowline where water ice unambiguously shows a higher threshold. There is evidently some uncertainty in the sticking efficiency of water ice aggregates, and it is common in models to consider the sticking and fragmentation thresholds for ice as parameters and explore the effect of different values in the 1-10 ms-1 range.
@@
Translating the experimental results for the sticking and fragmentation thresholds into a predicted maximum particle size $a_{\rm max}$ requires knowledge of what sets $\Delta v$ as a function of size $a$. In the most common situation, collisions at $a\sim a_{\rm max}$ are caused by aerodynamic coupling to small-scale turbulence (this process also leads to small-scale clustering of solids; Cuzzi et al. 2001). For particles of size $a_{1}$ and $a_{2}$, coupled aerodynamically to isotropic Kolmogorov turbulence, Ormel and Cuzzi (2007) provide the standard (astrophysical) calculation that gives $\Delta v(a_{1},a_{2})$. Adopting an $\alpha$ model for the turbulence, the peak collision speed is found to be,
@@
This allows observations of the thickness of dust disks to be used to constrain disk turbulence (Villenave et al. 2025). The assumption that $D=\nu$ is quite crude, and there can be substantial deviations from this simple formula in disks where the turbulence is driven by the Vertical Shear Instability (Stoll and Kley 2016; Pfeil et al. 2024; Fukuhara and Okuzumi 2024) or by self-gravity (Baehr and Zhu 2021).
@@
Even though $\eta\sim(h/r)^{2}$ is small, typically only $\sim{\rm few}\times 10^{-3}$, the rate of the resulting drift is highly consequential. At 1 au, the time scale for drift of particles with $\tau_{\rm S}=1$ is as short as of the order of a thousand years. Particles with this Stokes number would be of the order of meters in size, leading to the moniker “meter-sized barrier” for this aerodynamic constraint on planetesimal formation. (Material physics limits on sticking already make the growth of meter-sized boulders improbable, but the meter-sized barrier provides an independent, and arguably more robust, argument against planetesimal formation by pairwise coagulation.) More importantly in a modern context, particles with physical sizes of mm-cm and Stokes numbers $\tau_{\rm S}\sim 10^{-2}$, that can readily form given our best understanding of coagulation physics, are predicted to suffer substantial drift even in the outer disk on scales of 10-100 au.
@@
Turbulent diffusion in the radial direction can be important both for the evolution of trace gas species, and for the evolution of dust populations that are simultaneously subject to radial drift (Morfill and Voelk 1984; Clarke and Pringle 1988). Tominaga et al. (2019) derive formulae for dust diffusion in disks that conserve angular momentum, and which can be used to model dust concentration processes and possible instabilities. An important consequence of diffusion is the qualitative change it makes to how dust evolves in the vicinity of gas pressure maxima. Ignoring diffusion, an axisymmetric pressure maximum – whether it forms spontaneously or at a planetary gas edge – is efficient at trapping aerodynamically drifting solids, filtering all but the smallest particles that are swept inward by the mean flow (Rice et al. 2006). With diffusion, it is always possible to find a steady state solution in which the turbulent flux overcomes the aerodynamic trapping effect, such that pressure maxima temporarily detain but do not permanently imprison solids (Ward 2009). Numerical simulations confirm that so-called “particle traps”, are, generically, rather leaky (Zhu et al. 2012b; Huang et al. 2025).
@@ 4 Planetesimal formation
If $\rho_{\rm dust}>\rho_{\rm Roche}$, solid material can collapse gravitationally to form planetesimals. Although there are circumstances where this process may operate, it is no longer considered to be a generally viable planetesimal formation mechanism. Typically, vertical shear will generate enough turbulence to halt settling before the critical density is reached, even in an intrinsically non-turbulent disk (Cuzzi et al. 1993). The other obvious mechanism for planetesimal formation is a continuation of pairwise sticking collisions from pebble to planetesimal scales (Weidenschilling and Cuzzi 1993). While this model cannot be entirely ruled out, it is largely unviable given current understanding of fragmentation and radial drift processes.
The current working hypothesis for planetesimal formation is that the streaming instability (Youdin and Goodman 2005) generates particle clumps that are dense enough to collapse gravitationally to planetesimals. The streaming instability is a local, linear instability of the Nakagawa et al. (1986) solution for equilibrium radial drift. It requires consideration of the momentum feedback particles exert on the gas, and a non-zero drift velocity (hence, it cannot formally be operative at the peak of a pressure maximum). Youdin and Goodman (2005) show that the streaming instability has non-zero (though sometimes small) growth rates for almost all combinations of Stokes number and dust to gas ratio, and it should therefore be generically present in protoplanetary disks. Jacquet et al. (2011) and Magnan et al. (2024) discuss the rather subtle physical mechanism of the streaming instability, which is the most important (for planet formation) of a broader class of resonant drag instabilities (Squire and Hopkins 2018). It should be noted that the Nakagawa et al. (1986) equilibrium studied by Youdin and Goodman (2005) differs in multiple ways from both real disks and from the setups for most streaming instability simulations. The Nakagawa et al. (1986) model is vertically unstratified (whereas, in actual disks, vertical gravity is not normally negligible), does not include turbulent diffusion (which is present if there is any instrinsic disk turbulence and which is needed to establish a stratified particle layer equilibrium), and is based on a fluid description of the particles (this is only a good approximation for $\tau_{\rm S}\ll 1$, and is not used in particle-based simulations). Subsequent linear calculations have included the effects of a particle size distribution (Krapp et al. 2019; Paardekooper et al. 2020; Zhu and Yang 2021) and turbulent diffusion (Chen and Lin 2020; Umurhan et al. 2020; Gerbig et al. 2024). Both effects act to reduce the growth rates and restrict the range of parameters over which linear instability is present.
@@
A second important simulation result is the predicted mass function of planetesimals formed from the streaming instability. A large number of studies agree that streaming-initiated collapse leads to a top-heavy initial planetesimal mass function, in which most of the total mass is in the largest objects (Johansen et al. 2007; Simon et al. 2016; Schäfer et al. 2017; Li et al. 2019; Schäfer et al. 2024). The absolute size of the largest planetesimals is of course a function of the disk conditions, but for reasonable parameters it is often hundreds of kilometers. This first principles prediction is large compared to what was assumed in most formation-agnostic planetary growth models in the past, and has substantial downstream implications both for the Solar System (“asteroids were born big”; Morbidelli et al. 2009) and for planet formation generally. In particular, planetesimals that are 100 km and larger in size experience much less aerodynamic drag than 0.1-1 km planetesimals, and when their orbits are excited by gravitational perturbations the population will saturate at a higher velocity dispersion. This reduces the rate at which protoplanets grow by planetesimal accretion, so that, counter-intuitively, being able to rapidly make large primordial planetesimals does not necessarily imply faster overall planet formation time scales.
There is no analytic theory for the form of the initial planetesimal mass function, but several simple functions (an exponentially tapered power-law, a broken power-law, etc) provide fits to numerical results that can be compared to Solar System populations of small bodies. The Kuiper Belt is widely considered to be the most primordial surviving population and offers the greatest potential for a test, though even there careful consideration has to be given to the extent of collisional evolution (Morbidelli et al. 2021; Benavidez et al. 2022). Other possible tests of the predictions of planetesimal formation models include the properties and inclination distribution of Kuiper Belt binaries (Grundy et al. 2019; Porter et al. 2024), which are consistent with streaming instability expectations (Nesvorný et al. 2019). The detailed morphology of the single small Kuiper Belt Object imaged by the New Horizons mission (Arrokoth; McKinnon et al. 2020) also supports a streaming formation mechanism.
@@ 5 Formation of rocky planets, cores, and giants
    Planetesimal accretion. Planetesimals physically collide, due to gravitational interactions, at low enough velocity to grow. The rate of collisions is determined, through the physics of gravitational focusing, by the velocity dispersion of the planetesimals. To a first approximation, one imagines the protoplanet growing from a local planetesimal reservoir that does not drift radially.
@@
    Pebble accretion. While gas is still present, a fraction of the appropriately aerodynamically coupled pebbles (typically, those with $10^{-3}\leq\tau_{\rm S}\leq 1$) that enter the protoplanet’s Hill sphere are gravitationally captured and accreted via a settling flow. The protoplanet accretes a fraction of the solids that drift past its orbit, with an efficiency that depends on its mass, on the Stokes number, and on the disk properties.
@@
Applying the condition that the settling time scale $t_{\rm settle}=b/v_{\rm settle}$ equals the encounter time scale, $t_{\rm enc}=b/\delta v$, we find that $\delta v=GMt_{\rm stop}/b^{2}$, and the pebble accretion rate is,
@@
Pebble accretion has several distinctive differences from planetesimal accretion. There is no regime of pebble accretion that exhibits runaway growth. It is also fairly strictly bounded in planet mass. Pebble accretion is inefficient below a pebble initiation mass (Visser and Ormel 2016), and for typical parameters would only be able to directly grow planets from planetesimals if their initial sizes are rather large (at least several hundred km). Pebble accretion is also likely to be inefficient above a pebble isolation mass (Lambrechts et al. 2014), estimated by Bitsch et al. (2018) as,
$M_{\rm iso,pebble}\simeq 25\left(\frac{h/r}{0.05}\right)^{3}\left(\frac{M_{*}}{M_{\odot}}\right)fM_{\oplus},$ (20)
where $f$ is a correction factor that depends on the radial pressure gradient, viscosity, and strength of turbulent diffusion (see Chapter by Ormel in this volume). Note the strong dependence on $(h/r)$. Above this mass, a growing planet starts to open a gap in the disk, forming a local pressure maximum that can stop the radial drift of solids. Simulations show that this suppression of radial drift can be reduced by three dimensional effects (Huang et al. 2025).
@@
Planetesimal accretion is severely limited at large orbital radii by the tendency of more massive bodies to scatter, rather than accrete, planetesimals (this is the origin of the long-known difficulty in forming Uranus and Neptune in situ from planetesimals; Levison and Stewart 2001). Pebble accretion, conversely, tends to work better at moderately large orbital radii – solids at the drift or fragmentation limit have more favorable aerodynamic properties, and there is no reduction in accretion efficiency as the ratio of Hill sphere size to planetary radius increases. Outside perhaps 5-10 au, pebble accretion is presumptively the dominant mechanism for planetary growth (e.g. Drazkowska et al. 2023). More general statements as to the relative role of planetesimals and pebbles in planet formation are, however, hazardous. Planetesimal and pebble accretion are distinct processes that cannot be easily compared, and while estimated pebble accretion rates can be very high this is, in part, a consequence of the rapid radial drift of solids which cannot be sustained indefinitely. Which process dominates in the formation of observed planetary systems can only be considered in the context of a global theoretical model for planet formation (e.g. Bitsch et al. 2015, 2018), or by appealing to observational cosmochemical constraints that are only available for the Solar System. Johansen et al. (2021) argue for a large pebble contribution in terrestrial planet formation, but the division between the two processes in a hybrid Solar System model remains the subject of vigorous debate.
@@
Accretion of the primary atmospheres or envelopes of massive planets occurs (in the consensus view) via the process of core accretion (Mizuno 1980; Ikoma and Kobayashi 2025). Core accretion is a broad umbrella that covers the classical planetesimal-only models for Jupiter and Saturn’s formation (Pollack et al. 1996), along with more recent models that can accommodate the existence of the dilute core inferred for Jupiter from JUNO data (Wahl et al. 2017; Helled et al. 2022), or a primary role for pebble accretion (Lambrechts et al. 2014). The key insight underlying all these flavors of core accretion is the predicted existence of a sharp increase in the gas accretion rate when the envelope mass $M_{\rm env}$ exceeds a critical mass, that is approximately set at $M_{\rm env}\sim M_{\rm core}$ by the onset of self-gravity in the envelope (Rafikov 2006). Below the critical mass, low mass gaseous protoplanets grow relatively slowly, at a rate that is determined by the growth of the core through pebble and planetesimal accretion, and by the time scale on which the envelope can lose thermal energy and contract (Ikoma et al. 2000). If the gas disk dissipates during this initial phase of core accretion, the expected outcome is a planet that may resemble the Solar System’s ice giants, or extrasolar sub-Neptunes. If, instead, the critical mass can be exceeded during the disk lifetime, the planet enters a phase of accelerated gas accretion (this is known as “runaway growth”, although it is still a quasi-hydrostatic evolution) and the outcome is a gas giant with $M_{\rm env}\gg M_{\rm core}$. The minimum core mass that can lead to gas giant formation is primarily a function of the opacity in the envelope, and is roughly $5-10\ M_{\oplus}$ (Piso et al. 2015) for grain size distributions that arise in coagulation-fragmentation equilibrium calculations. Smaller critical core values are possible if envelopes are effectively grain-free (Hori and Ikoma 2010). In the other direction, substantially more massive cores can avoid entering runaway given sufficient heating from planetesimal accretion.
@@
[…]
The consensus view is that the Solar System’s planets, and at least the vast majority of extrasolar planets, formed via core accretion (which, as noted already, is today more of a broad guideline for planet formation rather than a precisely specified rule). It is possible that some giant planets, especially those approaching the $13\ M_{J}$ threshold that conventionally demarcates planets from brown dwarfs, instead formed via disk instability (Boss 1997). The formation of protostellar disks may well lead to initially self-gravitating disks (Lin and Pringle 1990; Xu and Kunz 2021), which have been inferred observationally from kinematic data (Speedie et al. 2024) and seen to fragment, at least into multiple stellar systems (Tobin et al. 2016). The physics of self-gravitating disks is reviewed by Kratter and Lodato (2016). Massive disks fragment if the cooling time scale is sufficiently short (of the order of the orbital time scale; Gammie 2001; Rice et al. 2005), conditions that occur only at large ($\sim 50$ au) radii (Clarke 2009; Rafikov 2009). What are currently the most realistic radiation hydrodynamics simulations, by Xu et al. (2024), find that the distribution of initial fragment masses in this regime peaks at about $20\ M_{J}$. Further accretion in what remains a gas-rich environment post-fragmentation would likely increase fragment masses further (Kratter et al. 2010b). (This fate could only be averted if fragments migrate inward rapidly and are partially tidally stripped; Nayakshin 2010; Baruteau et al. 2011; Zhu et al. 2012a). Protostellar disks can also be driven to fragment by sufficiently rapid infall (Kratter et al. 2010a), but this also leads to characteristic masses that are typically in the brown dwarf or low-mass star regime.